# HyDeFuse: Provably Convergent Denoiser-Driven Hyperspectral Fusion

### v1

@@ 3.3 Kernel Denoiser
CasKD $\mathbf{32.43}$ $\mathbf{0.955}$ $\mathbf{30.86}$ $\mathbf{0.931}$
Table 1: Comparison of denoising performance on Pavia and Paris. We see that CasKD generally does better than its components $\mathcal{V}$ and $\mathcal{W}$ and their other combinations. The top two methods are in bold and the best method is underlined.
@@
Figure 3: Visual comparison of CasKD with its component denoisers $\mathcal{V}$ and $\mathcal{W}$ for Gaussian noise with standard deviation $0.078$.
@@ 4 Convergence Analysis
In this section, we establish global linear convergence of (9) with (12) as the denoiser. Our analysis builds on the convergence framework presented in [49]. However, there are important technical differences that prevent the direct application of the results from [49]. These differences are best explained after presenting the technical results, so we postpone this discussion until the end of the section.
4.1 Fixed Point Iteration
The starting point in the analysis is to view (9) as a fixed point iteration. Specifically, we can write (9) as
@@ 4.1 Fixed Point Iteration
Thus, we can view (19) as a fixed point iteration
@@
We will refer to $\mathcal{G}$ as the gradient step operator and the affine map $\mathcal{T}$ as the fixed point operator.
We will later show that the operators $\mathcal{D}$ and $\mathcal{G}$ are nonexpansive. Consequently, $\mathcal{P}=\mathcal{D}\circ\mathcal{G}$, and hence $\mathcal{T}$, is nonexpansive. However, this is not sufficient to establish convergence of $\{\mathbf{X}_{k}\}$ generated by $\mathbf{X}_{k+1}=\mathcal{T}(\mathbf{X}_{k})$ for any arbitrary $\mathbf{X}_{0}\in\mathbb{H}$. A sufficient condition is that $\mathcal{P}$ (and hence $\mathcal{T}$) is a contraction operator, i.e., there exists $\mu\in[0,1)$ such that
$\|\mathcal{P}(\mathbf{X}_{1})-\mathcal{P}(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\mu\,\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$ (23)
If $\mathcal{T}$ is a contraction, we use the contraction mapping theorem [72] to conclude that the fixed point iteration $\mathbf{X}_{k+1}=\mathcal{T}(\mathbf{X}_{k})$ converges to a unique fixed point for any initialization $\mathbf{X}_{0}\in\mathbb{H}$. The smallest possible $\mu$ in (23) is called the contraction factor of $\mathcal{T}$.
The rest of this section is dedicated to proving that $\mathcal{T}$ is a contraction operator. As noted above, it suffices to show that $\mathcal{P}$ is a contraction. Moreover, since $\mathcal{P}$ is linear, (23) is equivalent to
@@
We have the following simple observation in this regard.
@@
Suppose $\mathbb{H}$ is a finite-dimensional vector space and $\mathcal{P}\in\mathcal{L}(\mathbb{H})$. Then $\mathcal{P}$ is a contraction if and only if $\|\mathcal{P}(\mathbf{X})\|<\|\mathbf{X}\|$ for all nonzero $\mathbf{X}\in\mathbb{H}$. In particular, the contraction factor $\mu$ in (23) is given by
$\mu=\max\Big{\{}\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}:\ \mathbf{X}\in\mathbb{H},\,\|\mathbf{X}\|_{\mathbb{H}}=1\Big{\}}.$ (25)
We remark that this need not hold if $\mathcal{P}$ is not linear. One direction of Proposition 2 is obvious. In the other direction, it suffices to show $\mu<1$. Now, the set of $\mathbf{X}\in\mathbb{H}$ such that $\|\mathbf{X}\|_{\mathbb{H}}=1$ forms a compact subset of $\mathbb{H}$ and the map $\mathbf{X}\mapsto\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}$ is continuous. Therefore, by the extreme value theorem, there exists $\mathbf{X}^{*}\in\mathbb{H}$ such that $\|\mathbf{X}^{*}\|_{\mathbb{H}}=1$ and $\mu=\|\mathcal{P}(\mathbf{X}^{*})\|_{\mathbb{H}}<\|\mathbf{X}^{*}\|_{\mathbb{H}}=1$.
@@ 4.2 Constituent Operators
To show that $\mathcal{P}$ is a contraction, we will establish some special properties of the underlying operators $\mathcal{D}$ and $\mathcal{G}$. Note that we can write (22) as
$\mathcal{G}(\mathbf{X})=\mathbf{X}-\gamma\mathcal{K}(\mathbf{X}),\quad\mathcal{K}(\mathbf{X}):=\mathbf{P}_{1}\mathbf{X}+\mathbf{X}\mathbf{P}_{2},$ (26)
where
$\mathbf{P}_{1}=\mathbf{A}^{\top}\!\mathbf{A}\quad\mbox{and}\quad\mathbf{P}_{2}=\lambda(\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top}.$ (27)
We claim that $\mathcal{K}$ is self-adjoint on $\mathbb{H}$. Indeed, for any $\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}$,
$\big{\langle}\mathcal{K}(\mathbf{X}_{1}),\mathbf{X}_{2}\big{\rangle}_{\mathbb{H}}$ $=$ $\mathrm{tr}\left((\mathbf{P}_{1}\mathbf{X}_{1}+\mathbf{X}_{1}\mathbf{P}_{2})^{\top}\mathbf{X}_{2}\right)$
where we have used the cyclic property of trace and the fact that $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are symmetric matrices.
Since $\mathcal{K}$ is self-adjoint, it follows from (26) that the gradient step operator $\mathcal{G}$ is self-adjoint. This implies that the eigenvalues of $\mathcal{G}$ are real. We will next show that if the step size $\gamma$ in (22) is sufficiently small, namely if
$0<\gamma<2/\beta,\quad\beta:=\sigma_{\max}(\mathbf{A})^{2}+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}.$ (28)
then $\mathcal{G}$ is nonexpansive. Since $\mathcal{G}$ is self-adjoint, this would imply that $\sigma(\mathcal{G})\subset[-1,1]$. Furthermore, we will show that $1\notin\sigma(\mathcal{G})$ if $\gamma$ is in the range given by (28). Specifically, we have the following result.
@@
For the step size choice in (28), $\mathcal{G}$ is nonexpansive and $\sigma(\mathcal{G})\subset(-1,1]$.
The nonexpansive of the gradient step operator is a standard result in convex optimization [72]. This relies on the Ballion-Hadded theorem [72], a fundamental result about differentiable convex functions that is applicable for our loss function (6). A function $\ell:\mathbb{H}\to\mathbb{R}$ is said to be $\beta$-smooth if it is differentiable and its gradient $\nabla\!\ell$ is $\beta$-Lipschitz, i.e.,
$\|\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})\|_{\mathbb{H}}\leqslant\beta\|\mathbf{X}_{1}-\mathbf{X}_{2}\|_{\mathbb{H}}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$ (29)
The Ballion-Hadded theorem tells us that if $\ell$ is convex and $\beta$-smooth, then
$\langle\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2}),\mathbf{X}_{1}-\mathbf{X}_{2}\ \rangle_{\mathbb{H}}\geqslant\frac{1}{\beta}\|\nabla\!\ell(\mathbf{X}_{1})-\nabla\!\ell(\mathbf{X}_{2})\|_{\mathbb{H}}^{2}\qquad(\mathbf{X}_{1},\mathbf{X}_{2}\in\mathbb{H}).$
It is clear that (6) is convex, being the squared norm of a linear function. Moreover, comparing (11) and (26), we have
Therefore (30) becomes
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\beta\|\mathbf{X}\|_{\mathbb{H}}\qquad(\mathbf{X}\in\mathbb{H}),$ (31)
and (29) becomes
We next estimate $\beta$. It follows from (26) and the triangle inequality, that
$\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}\leqslant\|\mathbf{P}_{1}\mathbf{X}\|_{\mathbb{H}}+\|\mathbf{X}\mathbf{P}_{2}\|_{\mathbb{H}}\leqslant\big{(}\sigma_{\max}(\mathbf{P}_{1})+\sigma_{\max}(\mathbf{P}_{2})\big{)}\|\mathbf{X}\|_{\mathbb{H}}.$
Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are given by (27), we can take $\beta$ in (31) and (32) to be
$\beta=\sigma_{\max}(\mathbf{A}^{\top}\!\mathbf{A})+\lambda\sigma_{\max}((\mathbf{E}\mathbf{R})(\mathbf{E}\mathbf{R})^{\top})=\sigma_{\max}(\mathbf{A})^{2}+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}.$ (33)
We can now easily show that $\mathcal{G}$ is nonexpansive. Note that
$\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}^{2}=\|\mathbf{X}-\gamma\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}=\|\mathbf{X}\|_{\mathbb{H}}^{2}+\gamma^{2}\|\mathcal{K}(\mathbf{X})\|_{\mathbb{H}}^{2}-2\gamma\langle\mathbf{X},\mathcal{K}(\mathbf{X})\ \rangle_{\mathbb{H}}.$ (34)
Thus, if $\gamma$ is in the range given by (28), we have from (32) and (34) that
To complete the proof of Proposition (3), we must show that 1-1 is not an eigenvalue of $\mathcal{G}$. Indeed, if $1\in\sigma(\mathcal{G})$, we would have
for some nonzero $\overline{\mathbf{X}}\in\mathbb{H}$. Thus, if $\gamma$ is in the range given by (28), we get from (31) that
However, this is not possible, implying that $1\notin\sigma(\mathcal{G})$.
Thus far, we focused on the gradient step operator, one of the two operators in (21). We next analyze the denoiser $\mathcal{D}$, which is defined in terms $\mathcal{W}$ and $\mathcal{V}$.
The properties of $\mathcal{W}$ and $\mathcal{V}$ are determined by that of $\mathbf{W}$ and $\mathbf{V}$; see (15) and (18). For instance, $\mathbf{W}$ and $\mathbf{V}$ are symmetric, and this forces $\mathcal{W}$ and $\mathcal{V}$ to be self-adjoint –– the choice of the inner product (4) is crucial in this regard. We will also require additional properties of $\mathcal{W}$ and $\mathcal{V}$ in our analysis, which are summarized in the following result.
@@
The operators $\mathcal{V}$ and $\mathcal{W}$ are self-adjoint and nonexpansive, $\sigma(\mathcal{W})\subset[0,1]$, and $\mathrm{fix}\,(\mathcal{W})$ is a subspace of $\mathbb{H}$ consisting of matrices $\mathbf{F}$ of the form
@@
Proposition 4 is a key result that relies on the intricate construction of the denoisers in Section 3.3. Indeed, it is a non-trivial task to construct a kernel-based denoiser with the above mathematical properties. As previously discussed, the construction builds on prior works [42, 52], and the core part of the proof is based on Proposition 1. We have worked out the details in Appendix 7.1.
We can use the observation that $\mathcal{V}$ is nonexpansive to simplify our analysis. Recall the definition of $\mathcal{D}$ and $\mathcal{P}$ in (12) and (21). Using the associativity of composition, we can write
$\mathcal{P}=(\mathcal{V}\circ\mathcal{W})\circ\mathcal{G}=\mathcal{V}\circ(\mathcal{W}\circ\mathcal{G}).$
Therefore, if we can establish that $\mathcal{W}\circ\mathcal{G}$ is a contraction, it will follow that $\mathcal{P}$ is a contraction. In other words, we can focus on the two operators $\mathcal{W}$ and $\mathcal{G}$.
@@
We know from Proposition 4 that the $\mathrm{fix}\,(\mathcal{W})$ consists of matrices $\mathbf{F}$ of the form (36). Thus, to establish the above result, it suffices to show that if $\mathbf{F}$ is of the form (36) and $\mathbf{F}\neq\mathbf{0}$, then $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$. Note that $\mathbf{F}\neq\mathbf{0}$ means that $c_{j}\neq 0$ for some $1\leqslant j\leqslant L_{s}$. We will show using contradiction that under this condition $\mathbf{F}\notin\mathrm{fix}\,(\mathcal{G})$. Suppose that $\mathcal{G}(\mathbf{F})=\mathbf{F}$. Then, from (26), we have $\mathcal{K}(\mathbf{F})=\mathbf{0}$, i.e.,
$\mathbf{P}_{1}\mathbf{F}+\mathbf{F}\mathbf{P}_{2}=\mathbf{0},$
Since $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ are symmetric and positive semidefinite, we know that the two terms on the right are nonnegative. Therefore, we must have
where $\|\cdot\|_{2}$ is the standard Euclidean norm on $\mathbb{R}^{N_{m}}$. In particular, since we assume that $c_{j}\neq 0$ for some $j$, we must have $\mathbf{A}\boldsymbol{e}=\mathbf{0}$. However, recall that $\mathbf{A}=\mathbf{S}\mathbf{B}$, where $\mathbf{S}$ and $\mathbf{B}$ are sampling and blur operators. Since $\mathbf{B}\boldsymbol{e}=\boldsymbol{e}$, i.e., blurring a constant-intensity image produces the same image, we get $\mathbf{0}=\mathbf{A}\boldsymbol{e}=\mathbf{S}(\mathbf{B}\boldsymbol{e})=\mathbf{S}\boldsymbol{e}$. This is the desired contradiction since we cannot get a null vector by sampling $\boldsymbol{e}$.
4.3 Main Result
It turns out that the condition $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})=\{\mathbf{0}\}$ is not only necessary but also sufficient to guarantee that $\mathcal{W}\circ\mathcal{G}$ is a contraction, but under some assumptions on $\mathcal{W}$ and $\mathcal{G}$. To show this, we need the following result.
@@ 4.3 Main Result
Suppose $\mathcal{T}\in\mathcal{L}(\mathbb{H})$ is self-adjoint and $\sigma(\mathcal{T})\subset(-1,1]$. Then, for any $\mathbf{X}\in\mathbb{H}$,
@@
The above result identifies a condition under which $\mathcal{T}$ acts as a contraction in the complement of the subspace $\mathrm{fix}\,(\mathcal{T})$. This is the closest an operator can come to being a contraction if it has a nontrivial $\mathrm{fix}\,(\mathcal{T})$, i.e., if $\mathrm{fix}\,(\mathcal{T})\neq\{\mathbf{0}\}$. Although the final conclusion in Lemma (1) appears in [49], we provide a proof in Appendix 7.2 for completeness. Importantly, we see from Propositions 3 and 4 that $\mathcal{W}$ and $\mathcal{G}$ satisfy the assumptions in Lemma 1.
We are now ready to prove our main result: the contractivity of $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})$ and, consequently, the global linear convergence of HyDeFuse.
@@
The fixed point operator $\mathcal{T}$ in (20) is a contraction. In particular, for any arbitrary initialization $\mathbf{X}_{0}$, the iterates of HyDeFuse generated by (9) converge linearly to the unique fixed point of $\mathcal{T}$.
We had earlier noted that $\mathcal{T}$ is a contraction if $\mathcal{W}\circ\mathcal{G}$ is a contraction. By Proposition 2, it suffices to show that for any $\mathbf{X}\neq\mathbf{0}$,
$\|(\mathcal{W}\circ\mathcal{G})(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}.$
There are two possibilities. Supposex $\mathbf{X}\in\mathrm{fix}\,(\mathcal{G})$. Then $(\mathcal{W}\circ\mathcal{G})(\mathbf{X})=\mathcal{W}(\mathbf{X})$. However, we also have from Proposition 5, that $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{W})$. Since $\mathcal{W}$ is self-adjoint and $\sigma(\mathcal{W})\subset(-1,1]$, we have from Lemma 1 that $\|\mathcal{W}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Therefore,
@@
Consider the other possibility $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{G})$. In this case, we have $\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Thus, since $\mathcal{W}$ is nonexpansive, we again have
@@
Thus, we have established that $\mathcal{W}\circ\mathcal{G}$ and hence $\mathcal{T}$ is a contraction.
The second part of Theorem 1 is a standard application of the contraction mapping theorem [72], which guarantees the existence of a fixed point $\mathbf{X}^{*}\in\mathrm{fix}\,(\mathcal{T})$. In particular, if $\mu\in[0,1)$ is the contraction factor of $\mathcal{T}$, then
@@
In particular, we have
$\|\mathbf{X}_{k}-\mathbf{X}^{*}\|_{\mathbb{H}}\leqslant\mu^{k}\|\mathbf{X}_{0}-\mathbf{X}^{*}\|_{\mathbb{H}}\qquad(k\geqslant 0).$ (37)
This completes the proof of Theorem 1, establishing the global linear convergence of the iterates of HyDeFuse.
The above analysis allows us to identify the step size $\gamma$ in (9) required to guarantee convergence. This in turn depends on $\beta$ given (33). Since $\mathbf{A}=\mathbf{S}\mathbf{B}$, we can be shown that $\sigma_{\max}(\mathbf{A})\leqslant 1$, giving us the bound $\beta\leqslant 1+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}$, where $\sigma_{\max}(\mathbf{E}\mathbf{R})$ can easily be estimated using SVD. We can obtain a more precise bound using the power method [67] on the linear map $\mathcal{K}$ in (26). Empirical results show that this is often much smaller than the above bound. However, the analytical bound is useful as it is easy to compute and ensures convergence.

### v2

@@ 3.2 Kernel Denoiser
$\mathcal{V}\circ\mathcal{W}$ (CasKD) $\mathbf{32.43}$ $\mathbf{0.955}$ $\mathbf{30.86}$ $\mathbf{0.931}$
Table 1: Comparison of denoising performance on Pavia and Paris. CasKD generally performs better than its individual components, $\mathcal{V}$ and $\mathcal{W}$, as well as other combinations, except for $\mathcal{W}\circ\mathcal{V}$, for which its performance is comparable. The top two methods are shown in bold, and the best-performing method is underlined.
The CasKD module has several hyperparameters that affect the denoising process. Parameters such as the number of clusters and the window size are kept fixed in most experiments (typically $40$ and $10\times 10$), and the performance is generally robust to moderate changes in these values. However, other parameters—such as the patch size and the denoiser standard deviations $(\sigma_{1},\sigma_{2})$—are more problem-dependent. For instance, at higher noise levels, larger values of $(\sigma_{1},\sigma_{2})$ improve robustness, while increasing the patch size tends to reduce smoothing. Because the denoiser directly affects the reconstruction step in the fusion process, these parameters also influence the final image quality. In our experiments, we set $\sigma_{1}$ and $\sigma_{2}$ within the range $(5/255,25/255)$ and fixed the patch size at $7\times 7$. We observed that varying these parameters can lead to differences of about 4–5 dB in PSNR and 0.03–0.04 in UIQI, indicating their strong impact on fusion quality. Therefore, selecting appropriate values for $\sigma_{1}$, $\sigma_{2}$, and the patch size is crucial for achieving optimal performance.
@@
Figure 3: A visual comparison of CasKD on the Pavia dataset and its component denoisers, $\mathcal{V}$ and $\mathcal{W}$, in the presence of Gaussian noise with a standard deviation of $20/255$.
@@ 4 Convergence Analysis
We establish global linear convergence of HyDeFuse in this section. Our analysis builds on the convergence framework of [62]. However, there are key technical differences that prevent a direct application of those results. This will be clarified in Section 4.4, after we present our results. Proofs of all propositions and lemmas are given in the Appendix.
4.1 Fixed-Point Analysis
The starting point of our analysis is to interpret (9) as a fixed-point iteration. Specifically, we can rewrite (9) as
@@ 4.1 Fixed-Point Analysis
Thus, (20) can be expressed as
@@
We will refer to $\mathcal{G}$ as the gradient-step operator and the operator $\mathcal{T}$ as the fixed-point operator.
As we will see, both $\mathcal{D}$ and $\mathcal{G}$ are nonexpansive. Hence their composition $\mathcal{P}=\mathcal{D}\circ\mathcal{G}$, and consequently $\mathcal{T}$, is also nonexpansive. However, nonexpansivity alone does not guarantee convergence of the sequence $\{\mathbf{X}_{k}\}$. A sufficient condition for convergence is that $\mathcal{P}$ (and hence $\mathcal{T}$) is a contraction. In this case, the contraction mapping theorem [74] ensures that the sequence $\{\mathbf{X}_{k}\}$ converges to a unique fixed point for any initialization $\mathbf{X}_{0}$.
In the rest of this section, we prove that $\mathcal{P}$ is a contraction operator. Since $\mathcal{P}\in\mathcal{L}(\mathbb{H})$, it is a contraction if and only if there exists $\mu\in[0,1)$ such that
@@
We have the following observation in this regard.
@@
The operator $\mathcal{P}$ is a contraction if and only if $\|\mathcal{P}(\mathbf{X})\|<\|\mathbf{X}\|$ for all nonzero $\mathbf{X}\in\mathbb{H}$. In this case, we can take
$\mu=\max\big\{\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}:\ \mathbf{X}\in\mathbb{H},\,\|\mathbf{X}\|_{\mathbb{H}}=1\big\}$ (25)
as the contraction factor.
One direction is obvious. For the reverse direction, assume that $\|\mathcal{P}(\mathbf{X})\|<\|\mathbf{X}\|$ for all nonzero $\mathbf{X}\in\mathbb{H}$. Note that with $\mu$ defined in (25), we have (24); we just have to show that $\mu<1$. Since $\mathbb{H}$ is a finite-dimensional vector space, the set $\mathbb{S}=\{\mathbf{X}\in\mathbb{H}:\,\|\mathbf{X}\|_{\mathbb{H}}=1\}$ is compact. Moreover, the map $\mathbf{X}\mapsto\|\mathcal{P}(\mathbf{X})\|_{\mathbb{H}}$ is continuous on $\mathbb{S}$. Therefore, by the extreme value theorem, there exists $\mathbf{X}^{*}\in\mathbb{S}$ such that the maximum in (25) is attained, i.e., $\mu=\|\mathcal{P}(\mathbf{X}^{*})\|_{\mathbb{H}}$. It follows that $\mu=\|\mathcal{P}(\mathbf{X}^{*})\|_{\mathbb{H}}<\|\mathbf{X}^{*}\|_{\mathbb{H}}=1$.
@@ 4.2 Constituent Operators
To show that $\mathcal{P}$ is a contraction, we need some properties of $\mathcal{W},\mathcal{V}$ and $\mathcal{G}$. Note that we can express (23) as
$\mathcal{G}(\mathbf{X})=\mathbf{X}-\gamma\mathcal{K}(\mathbf{X}),$ (26)
where $\mathcal{K}$ is given by (12). We record an important property of $\mathcal{K}$.
@@
The operator $\mathcal{K}$ is self-adjoint and $\beta$-Lipschitz with $\beta=\sigma_{\max}(\mathbf{A})^{2}+\lambda\,\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}$.
Since $\mathbf{A}=\mathbf{S}\mathbf{B}$, it can be shown that $\sigma_{\max}(\mathbf{A})\leqslant 1$. This results in the bound $\beta\leqslant 1+\lambda\sigma_{\max}(\mathbf{E}\mathbf{R})^{2}$, where $\sigma_{\max}(\mathbf{E}\mathbf{R})$ can be easily estimated using SVD. A tighter bound can be obtained using the power method [70] applied to the linear map $\mathcal{K}$ in (26). Empirical results show that this estimate is often much smaller than the analytical upper bound, although the latter remains useful due to its simplicity and ease of computation.
As $\mathcal{K}$ is self-adjoint, it follows from (26) that $\mathcal{G}$ is also self-adjoint. Hence, all eigenvalues of $\mathcal{G}$ are real. We now show that if the step size $\gamma$ in (23) is chosen sufficiently small, then $\mathcal{G}$ is nonexpansive. Because $\mathcal{G}$ is self-adjoint, this further implies that $\sigma(\mathcal{G})\subset[-1,1]$. Moreover, we will establish that $1\notin\sigma(\mathcal{G})$. The result is summarized in the following proposition.
@@
For any $0<\gamma<2/\beta$, the operator $\mathcal{G}$ is nonexpansive and $\sigma(\mathcal{G})\subset(-1,1]$.
We note that the nonexpansiveness of the gradient-step operator is a standard result in convex optimization [74]. This follows from the Baillon–Haddad theorem [74, 75], a fundamental result for smooth convex functions that applies to our loss function(6). The statement about the spectrum, however, is specific to our operator.
Thus far, we focused on the gradient-step operator $\mathcal{G}$, one of the two operators in (22). We next analyze the denoiser $\mathcal{D}$, which is defined in terms $\mathcal{W}$ and $\mathcal{V}$.
The properties of $\mathcal{W}$ and $\mathcal{V}$ are determined by that of $\mathbf{W}$ and $\mathbf{W}^{(b)}$; see (16) and (19). For instance, $\mathbf{W}$ and $\mathbf{W}^{(b)}$ are symmetric, which forces $\mathcal{W}$ and $\mathcal{V}$ to be self-adjoint. The choice of the inner product (4) is crucial in this regard. We will also require additional properties of $\mathcal{W}$ and $\mathcal{V}$ in our analysis, which are summarized in the following result.
Proposition 5.
The operators $\mathcal{V}$ and $\mathcal{W}$ are self-adjoint and nonexpansive, $\sigma(\mathcal{W})\subset[0,1]$, and $\mathrm{fix}\,(\mathcal{W})$ is a subspace of $\mathbb{H}$ consisting of matrices of the form
@@
Proposition 5 is a key result that stems from the design of $\mathcal{W}$ and $\mathcal{V}$ in Section 3.2. Constructing a kernel denoiser with the stated mathematical properties is a nontrivial task.
We can use the fact that $\mathcal{V}$ is nonexpansive to simplify the analysis. From (13)and (22), we have
$\mathcal{P}=(\mathcal{V}\circ\mathcal{W})\circ\mathcal{G}=\mathcal{V}\circ(\mathcal{W}\circ\mathcal{G}).$ (28)
Hence, if we can show that $\mathcal{W}\circ\mathcal{G}$ is a contraction, it immediately follows that $\mathcal{P}$ is also a contraction. In other words, the analysis can be restricted to the two operators $\mathcal{W}$ and $\mathcal{G}$.
4.3 Linear Convergence
@@
It turns out that, under suitable assumptions on $\mathcal{W}$ and $\mathcal{G}$, the condition $\mathrm{fix}\,(\mathcal{W})\cap\mathrm{fix}\,(\mathcal{G})={\mathbf{0}}$ is not only necessary but also sufficient to ensure that $\mathcal{W}\circ\mathcal{G}$ is a contraction. To establish this, we first state the following result.
@@
Suppose $\mathcal{T}\in\mathcal{L}(\mathbb{H})$ is self-adjoint and $\sigma(\mathcal{T})\subset(-1,1]$. Then for any $\mathbf{X}\in\mathbb{H}$,
@@
The above result identifies a condition under which $\mathcal{T}$ acts as a contraction in the complement of the subspace $\mathrm{fix}\,(\mathcal{T})$. This is the closest a linear operator can come to being a contraction if it has a nontrivial $\mathrm{fix}\,(\mathcal{T})$, i.e., if $\mathrm{fix}\,(\mathcal{T})\neq\{\mathbf{0}\}$. Although a similar conclusion appears in [62], we include a proof in Appendix 7.5 for completeness. Importantly, Propositions 4 and 5 confirm that $\mathcal{W}$ and $\mathcal{G}$ satisfy the assumptions of Lemma 1.
We are now ready to establish our main result, namely that $\mathcal{T}$ is a contraction operator.
@@
The fixed-point operator $\mathcal{T}$ is a contraction. Consequently, for any initialization $\mathbf{X}_{0}\in\mathbb{H}$, the iterates generated by (9) converge linearly to the unique fixed point of $\mathcal{T}$.
Proof.
From (28), $\mathcal{P}$, and consequently $\mathcal{T}$, is a contraction if $\mathcal{W}\circ\mathcal{G}$ is a contraction. Furthermore, by Proposition 2, it suffices to show that for any $\mathbf{X}\neq\mathbf{0}$,
$\|(\mathcal{W}\circ\mathcal{G})(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}.$ (29)
There are two possibilities. Suppose $\mathbf{X}\in\mathrm{fix}\,(\mathcal{G})$. In this case, $(\mathcal{W}\circ\mathcal{G})(\mathbf{X})=\mathcal{W}(\mathbf{X})$. From Proposition 6, we know that $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{W})$. Since $\mathcal{W}$ is self-adjoint and $\sigma(\mathcal{W})\subset(-1,1]$, Lemma 1 implies that $\|\mathcal{W}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Hence,
@@
Consider the other possibility $\mathbf{X}\notin\mathrm{fix}\,(\mathcal{G})$. Here, we have $\|\mathcal{G}(\mathbf{X})\|_{\mathbb{H}}<\|\mathbf{X}\|_{\mathbb{H}}$. Thus, since $\mathcal{W}$ is nonexpansive,
@@
Thus, we obtain (29) in both cases. It follows that $\mathcal{W}\circ\mathcal{G}$, and consequently $\mathcal{T}$, is a contraction.
By the contraction mapping theorem [74], there exists a unique $\mathbf{X}^{*}\in\mathrm{fix}\,(\mathcal{T})$. If $\mu\in[0,1)$ denotes the contraction factor of $\mathcal{T}$, then
@@
By iterating this inequality, we obtain for all $k\geqslant 0$,
$\|\mathbf{X}_{k}-\mathbf{X}^{*}\|_{\mathbb{H}}\leqslant\mu^{k}\|\mathbf{X}_{0}-\mathbf{X}^{*}\|_{\mathbb{H}},$ (30)
which establishes the global linear convergence of the sequence $\{\mathbf{X}_{k}\}$. ∎
From (30), it follows that the convergence rate is directly governed by the contraction factor $\mu$. A smaller $\mu$ yields faster convergence of the iterates, whereas values of $\mu$ closer to $1$ lead to slower, yet still guaranteed, convergence.